Tell whether each number is divisible by 2,3,4,5,6,9,10. Then classify the number as even or odd?
1.(1,000), 2.(6,598), 3.(399), 4.(27,453), 5.(33,324) and 6.(16,335)?

Answer by AnlytcPhil(1276)

Tell whether each number is divisible by 2,3,4,5,6,9,10.  Then classify 
the number as even or odd?

1.(1,000), 2.(6,598), 3.(399), 4.(27,453), 5.(33,324) and 6.(16,335)?

-----------------------------------

First, I'll give you the rules and an example of each.

You'll need this: 0 is ALWAYS divisible by any positive integer.

A number is divisible by 2 if the last digit is 0, 2, 4, 6 or 8.
Example: 168 is divisible by 2 since the last digit is 8. 

A number is divisible by 3 if the sum of the digits is divisible by 3. 
Example: 168 is divisible by 3 since the sum of the digits is 15 (1+6+8=15),
and 15 is divisible by 3. 

A number is divisible by 4 if the number formed by the last two digits is
divisible by 4. 
Example: 316 is divisible by 4 since 16 is divisible by 4. 

A number is divisible by 5 if the last digit is either 0 or 5. 
Example: 195 is divisible by 5 since the last digit is 5. 

A number is divisible by 6 if it is divisible by 2 AND it is also divisible 
by 3. 
Example: 168 is divisible by 6 since it is divisible by 2 AND it is also
divisible by 3. 

A number is divisible by 8 if the number formed by the last three digits is
divisible by 8.
Example: 7,120 is divisible by 8 since 120 is divisible by 8. 

A number is divisible by 9 if the sum of the digits is divisible by 9.
Example: 549 is divisible by 9 since the sum of the digits is 18 (5+4+9=18),
and 18 is divisible by 9. 

A number is divisible by 10 if the last digit is 0.
Example: 1,470 is divisible by 10 since the last digit is 0. 

--------------------------------------

1.(1,000), 

1000 is divisible by 2 since it ends in 0

1000 is not divisible by 3 since the sum of the digits is 1 (1+0+0+0=1),
and 1 is not divisible by 3. 

1000 is divisible by 4 since 00, same as 0, is divisible by 4.

1000 is divisible by 5 since the last digit is 0.

1000 is not divisible by 6 because even though it is divisible by 2
it is not also divisible by 3.

1000 is divisible by 8 since 000, same as 0, is divisible by 8.

1000 is not divisible by 9 since the sum of the digits is 1 (1+0+0+0=1),
and 1 is not divisible by 9.

1000 is divisible by 10 since its last digit is 0.

-------------------------------------------------

2.(6,598),

6,598 is divisible by 2 since it ends in 8

6,598 is not divisible by 3 since the sum of the digits is 1 (6+5+9+8=28),
and 28 is not divisible by 3. 

6,598 is not divisible by 4 since 98 is not divisible by 4.

6,598 is not divisible by 5 since the last digit is not 0 or 5.

6,598 is not divisible by 6 because even though it is divisible by 2
it is not also divisible by 3.

6,598 is not divisible by 8 since 598 is not divisible by 8.

6,598 is not divisible by 9 since the sum of the digits is 28 (6+5+9+8=28),
and 28 is not divisible by 9.

6,598 is not divisible by 10 since its last digit is not 0.

-------------------------------------------------------------

You do the rest.  Just follow the rules above:

 3.(399)

Answer: divisible by 3 only


 4.(27,453)

Answer: divisible by 3 only 


 5.(33,324) 

Answer: divisible by 2, 3, 4, 6 

 6.(16,335)

Answer: divisible by 3, 5, 9

Edwin